On the zeros of q-Laguerre polynomials

Preeti Agarwal Motilal Nehru National Institute of Technology, Allahabad, India

q - laguerre polynomials are an orthogonal polynomial system for α > −1.In this paper the behavior of zeros in the neighborhood of x = 0 is considered for α − k, where k = 2. It is shown that limit of the sum of the reciprocals of the zeros as α → −k is finite.